40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 14 23:56:18 2019, rana, Update, LSC, off the bad Violin filters out.pdf
    Reply  Fri Nov 15 11:58:12 2019, gautam, Update, LSC, off the bad Violin filters violinFix.pdfnewViolinConfig.png
    Reply  Mon Mar 1 16:27:17 2021, gautam, Update, LSC, PRM violin filter excessive? violins.pdfPRviolin.pdf
       Reply  Tue Mar 2 13:39:31 2021, rana, Update, LSC, PRM violin filter excessive? 
Message ID: 15028     Entry time: Fri Nov 15 11:58:12 2019     In reply to: 15026
Author: gautam 
Type: Update 
Category: LSC 
Subject: off the bad Violin filters 

The clue was that the loop X arm POX loop looked to have low (<3dB)) gain margin around 600 Hz (and again at 700 Hz). Attachment #1 shows a comparison of the OLTF for this loop (measured using the IN1/IN2 method) before and after our change. We hypothesize that one of the violin filters that were turned off had non-unity DC gain, because I had to lower the loop gain by 20% after these turn-offs to have the same UGF. I updated the snap files called by the arm locking scripts.

I think I caught all the places where the FM settings are saved, but some locking scripts may still try and turn on some of these filters, so let's keep an eye on it.

Quote:

We turned off many excessive violin mode bandstop filters in the LSC.

Attachment 1: violinFix.pdf  10 kB  | Show | Hide all | Show all
Attachment 2: newViolinConfig.png  28 kB  | Hide | Hide all | Show all
newViolinConfig.png
ELOG V3.1.3-